Q1. [40 marks] Answer the following multiple-choice questions (Please note that to score full marks in a question, you have t l
Answer #1

We have the following information

Demand for bread = 20,000 loaves per year

Per loaf cost = $0.50

Annual cost of purchasing the loaf = 20,000 × 0.50 = $10,000

Machine A

Machine B

Capital investment ($)

8,000

12,000

Useful life (n) in years

7

7

Annual fixed cost ($)

2,000

3,500

MARR (i)

10% or 0.1

10% or 0.1

Machine A

Equivalent uniform annual cost (EUAC) = First Cost(A/P, i, n) + Annual Fixed Cost

EUAC = 8000(A/P, 10%, 7) + 2000

EUAC = 8000[0.1(1 + 0.1)7/((1 + 0.1)7 – 1)] + 2000

EUAC = 1643.24 + 2000

EUAC of Machine A = $3,643.24

Machine B

EUAC = First Cost(A/P, i, n) + Annual Fixed Cost

EUAC = 12000(A/P, 10%, 7) + 3500

EUAC = 12000[0.1(1 + 0.1)7/((1 + 0.1)7 – 1)] + 3500

EUAC = 2464.84 + 3500

EUAC of Machine B = $5,964.84

The EUAC of both Machine A and Machine B is less than the annual cost of buying bread. In fact, the annual cost of buying bread is higher than the combined EUAC of both Machine A and Machine B. So, the supermarket chain can buy both the machines.